# Estimation of Optimal Causal Bounds via Covariate-Assisted Optimal Transport

### v1

@@ 4.2 Covariate-Dependent Treatment Assignment
Unlike Bernoulli design where treatment assignment is independent of covariates by design, observational studies typically exhibit covariate-dependent treatment assignment. This results in non-constant propensity scores
$e(z)=\mathbb{P}(W_{i}=1\mid Z_{i}=z),$
reflecting heterogeneity in treatment likelihood across covariate profiles. In this section, we discuss how to modify the adapted COT value estimator to fit this situation and characterize the associated finite-sample convergence rate.
@@
Under the unconfoundedness assumption (Assp. 2.2), the conditional distributions of potential outcomes given covariates remain identifiable, i.e., consistent with $\mu_{Y(k)}^{z}$ for $k=0,1$. However, since the treatment assignment probability depends on covariates, the marginal distribution of covariates may differ between the treatment group and the control group, inducing a covariate shift. To formalize this setting with covariate shift, we consider the following super-population model of $Y(0),Y(1),Z,W$ following the joint distribution $\mu=\mu_{Y(0),Y(1),Z,W}$.
Assumption 4.3 (Covariate shift).
Assume that $((Z_{i}(0),Y_{i}(0)),i\in[n])$ are drawn i.i.d. from $\mu_{Z|W=0}\otimes\mu_{Y(0)}^{Z}$, and $((Z_{j}(1),Y_{j}(1)),j\in[m])$ are drawn i.i.d. from $\mu_{Z|W=1}\otimes\mu_{Y(1)}^{Z}$.
@@
Since our estimand of interest $V^{*}$ relies on the marginal covariate distribution, irrespective of treatment assignment, it is essential to correct for the distributional shift induced by non-constant propensity scores. Specifically, we reweight the samples to align with the marginal covariate distribution. By Proposition 2.1, when the propensity score $e(z)$ is known, the appropriate population-level weighting is given by
$w_{i}=\frac{\mu_{Z}(z)}{\mu_{Z\mid W=0}(z)}\propto(1-e(z))^{-1},\ W_{i}=1,\quad\xi_{j}=\frac{\mu_{Z}(z)}{\mu_{Z\mid W=1}(z)}\propto e^{-1}(z),\ W_{j}=0.$
@@
On finite samples, when the propensity score $e(z)$ is unknown but estimated by $\hat{e}(z)$, we adopt the self-normalized weight,
$\hat{w}_{i}=\frac{(1-\hat{e}(Z_{i}(0)))^{-1}}{\sum_{l=1}^{n}(1-\hat{e}(Z_{l}(0)))^{-1}},\ i\in[n],\quad\hat{\xi}_{j}=\frac{\hat{e}^{-1}(Z_{j}(1))}{\sum_{l=1}^{m}\hat{e}^{-1}(Z_{l}(1))},\ j\in[m].$ (9)
@@
to avoid extreme values and ensure the weights sum to one. To adapt the construction of the adapted COT value estimator with discretization to the setting in Assp. 4.3, we modify the weights in Eq. (9) by evaluating the estimated propensity score at the projected covariate values rather than the original values. Finally, denoting the cells corresponding to projection map $\varphi_{r}^{n}$ by $\Phi_{r}^{n}\triangleq\left\{(\varphi_{r}^{n})^{-1}(\{x\}):x\in\varphi_{r}^{n}([0,1]^{d_{Z}})\right\}$, our reweighted adapted empirical distributions are defined as
$\hat{\bm{\mu}}_{Y(0),Z(0),n}$ $=\sum_{G\in\Phi_{r}^{n}}\sum_{i:Z_{i}(0)\in G}\frac{(1-\hat{e}(\varphi_{r}^{n}(G)))^{-1}}{\sum_{G}(1-\hat{e}(\varphi_{r}^{n}(G)))^{-1}|\{i:Z_{i}(0)\in G\}|}\delta_{Y_{i}(0),\varphi_{r}^{n}(Z_{i}(0))},$
$\hat{\bm{\mu}}_{Y(1),Z(1),m}$ $=\sum_{G\in\Phi_{r}^{m}}\sum_{j:Z_{j}(1)\in G}\frac{\hat{e}(\varphi_{r}^{m}(G))^{-1}}{\sum_{G}\hat{e}(\varphi_{r}^{m}(G))^{-1}|\{j:Z_{j}(1)\in G\}|}\delta_{Y_{j}(1),\varphi_{r}^{m}(Z_{j}(0))},$
@@
where $\varphi_{r}^{n}(G)$ is the cell center of $G\in\Phi_{r}^{n}$. We further employ $\hat{\bm{\mu}}_{Y(0),Z(0),n}$ and $\hat{\bm{\mu}}_{Y(1),Z(1),m}$, along with $\hat{\pi}_{n,m}$ defined in Eq. (8), to construct the adapted COT value estimator introduced in Definition 4.3.
@@
We implement a two-fold cross-fitting procedure to ensure the independence between the estimated propensity score $\hat{e}$ and the adapted empirical distributions $\hat{\bm{\mu}}$. Specifically, we randomly split the data into two folds of approximately equal size. For each fold, we construct $\hat{e}$ using data from the other fold, and then construct $\hat{\bm{\mu}}_{Y(0),Z(0),n},\hat{\bm{\mu}}_{Y(1),Z(1),m}$ on the current fold with the estimated propensity score plugged in. This yields two estimates of the lower bound, and we further average them to form our final estimator $\hat{V}_{n,m}$.
We introduce the assumptions on the estimation error of $\hat{e}(\cdot)$, which are used by the convergence analysis of $\hat{V}_{n,m}$.
Assumption 4.4 (Lipschitz $\hat{e}$).
@@
Assumption 4.6 (Average error of propensity score).
Assume that for $m,n\geq 1$,
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left|n\hat{w}_{i}-\frac{{\rm d}\mu}{{\rm d}\mu_{|W=0}}(Z_{i})\right|\right]\leq C_{w}n^{-r},\quad\mathbb{E}\left[\frac{1}{m}\sum_{j=1}^{m}\left|m\hat{\xi}_{i}-\frac{{\rm d}\mu}{{\rm d}\mu_{|W=1}}(Z_{i})\right|\right]\leq C_{w}m^{-r},$
where $C_{w}$ is a constant independent of $n,m$.
Assp. 4.6 characterizes the average estimation error of the estimated propensity score function $\hat{e}$. Notably, we impose a mild convergence rate requirement of order $O(n^{-r})$, where $r$ is chosen to be less than $1/d_{Z}$. This rate aligns with typical nonparametric convergence rates for nuisance parameter estimation. Assp. 4.4 and 4.5 impose boundedness and smoothness conditions on $\hat{e}(\cdot)$. Particularly, the smoothness condition is necessary because we evaluate $\hat{e}$ at the centers of discretized cells, as discussed in the following.
Theorem 4.3 (Finite-sample complexity with covariate shift).
Under Assp. 2.33.1 -  3.4,  4.3 - 4.6 with $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, we have
$\mathbb{E}\left[|\hat{V}_{n,m}-V_{\textup{c}}|\right]\leq C^{\dagger}\gamma_{d_{Y},d_{Z}}(m\wedge n),$
where $C^{\dagger}$ is a constant that depends on $d_{Z},d_{Y}$, $\delta$ (Assp. 2.3), $M$ (Assp. 3.2), $L_{h}$ (Assp. 3.3), $L_{Z}$ (Assp. 3.4), $L_{e}$ (Assp. 4.4), $\eta$ (Assp. 4.5) and $C_{w}$ (Assp. 4.6).
The error rate of $\hat{V}_{n,m}$ under non-constant propensity scores (Theorem 4.3) matches that in the case of homogeneous propensity scores (Theorem 4.2).
@@ 5 Simulation
The code can be found in the link: https://github.com/siruilin1998/COT_estimation.git.
5.1 Synthetic Data
5.1.1 Bernoulli Design
To evaluate the effectiveness of the adapted COT value estimator in simulation [^5] All code and additional experimental results are provided in the supplementary material., we introduce the following synthetic data-generating mechanism. For $k=0,1$, $Y(k)=V_{m}(f_{k}(Z),\varepsilon_{k})$, where $Z,\varepsilon_{k},Y(k)$ are independent with each other. The model $V_{m}$ includes: (i) location model $V_{1}(u,v)=u+v$, and (ii) scale model $V_{2}(u,v)=(u_{j}v_{j},j\in[d_{Y}])$.
More concretely, in this section, we consider $d_{Y}=d_{Z}=1$, $Z,\varepsilon_{0},\varepsilon_{1}\sim N(0,1)$, and three models:
@@ 5.1.1 Bernoulli Design
For Gaussian models, by the following result, the true value of $V_{\textup{c}}$ is computable through Monte Carlo algorithm with parametric convergence rate, and can serve as the true value for evaluating the experimental results.
Lemma 5.1 ($V_{\textup{c}}$ for Gaussian noise model, [lin2025tightening, Lemma 6.1]).
  1. (i)
  2. (ii)
@@
We compare the estimation accuracy of our adapted COT value estimator with the indirect method proposed by [ji2023model], implemented in their Python package DualBounds [^6] https://dualbounds.readthedocs.io. Their approach relies on estimating nuisance functions associated with the covariate–outcome model. In our comparison, we evaluate both the ridge regression-based and $k$-nearest neighbor (KNN)-based DualBounds methods. Our implementation is based on the Python Optimal Transport (POT [^7] https://pythonot.github.io/) library [flamary2021pot], which we use to solve the OT problems involved in our method. Experiments were run on a MacBook Air (Apple M3, 8-core CPU, 16 GB RAM), without GPU acceleration.
(a) Linear location model
   (b) Quadratic location model
 (c) Scale model
Figure 2: Estimation accuracy comparison between the adapted COT value estimator and DualBounds. The average relative error is defined as $\mathbb{E}\left[\left|\text{estimator}-V_{\textup{c}}\right|\right]/V_{\textup{c}}$, where the expectation $\mathbb{E}$ is approximated by averaging over 500 Monte Carlo repetitions. To quantify the uncertainty, we compute the standard error of the mean (SEM) as $\textup{SD}/\sqrt{500}$, where SD denotes the standard deviation of the relative errors across repetitions. We report the numerical values of SEM below to provide a measure of estimation variability. The maximum SEM values are: less than 0.005 in Figure 2(a), less than 0.02 in Figures 2(b) and 2(c). In the legend, “adapt” refers to the adapted COT value estimator, “ridge” refers to the ridge regression–based DualBounds method, and “knn” refers to the KNN–based DualBounds method.
@@
The results are presented in Figure 2. We can see that: (a) For the linear location model, the ridge-based DualBounds method performs best due to its use of a linear nuisance estimator that aligns with the true model structure. Nonetheless, the adapted COT value estimator achieves comparable accuracy despite not relying on model-specific assumptions. As the sample size increases, the performance gap quickly narrows. (b) For the quadratic location model, the adapted COT value estimator significantly outperforms both versions of DualBounds. In this case, the KNN-based DualBounds method outperforms the ridge-based version, as the latter cannot capture the nonlinearity in the true model. (c) For the scale model, the adapted COT value estimator consistently achieves the best performance. Notably, the KNN-based DualBounds method can produce negative estimates despite the true COT value being positive, highlighting the risk of using an ill-suited nuisance estimator.
Furthermore, we consider a setting in which the covariates of the treatment group are perturbed. Specifically, we set $\mu_{Z}=\mu_{Z(0)}=\mathcal{N}(0,1)$ and perturb the treatment covariate distribution as $\mu_{Z(1)}=\mathcal{N}(0,1)+\eta\varepsilon$, where $\varepsilon$ is an independent $\mathcal{N}(0,1)$ noise and $\eta$ controls the magnitude of the perturbation. Using the same three models described above, we evaluate the estimation accuracy of the adapted COT value estimator under varying levels of $\eta$. The results are reported in Figure 3. We observe that the adapted COT value estimator remains stable across small perturbation levels ($\eta\in[0,0.1]$), illustrating its robustness to model misspecification arising from covariate distribution mismatch (Remark 4.1). Intuitively, this robustness partially stems from the projection of covariates onto a finite set of representative values: small perturbations in the covariates are often absorbed during coarsening, as they rarely alter cell assignments. Moreover, the use of optimal coupling to match the discretized marginals helps mitigate the impact of systematic perturbations such as location shifts.
5.1.2 Covariate-Dependent Treatment Assignment
(a) Linear location model
 (b) Quadratic location model
(c) Scale model
Figure 4: Estimation accuracy comparison between the adapted COT value estimator and DualBounds. The average relative error is defined as $\mathbb{E}\left[\left|\text{estimator}-V_{\textup{c}}\right|\right]/V_{\textup{c}}$, where the expectation $\mathbb{E}$ is approximated by averaging over 500 Monte Carlo repetitions. We report the numerical values of standard error of the mean (SEM) below to provide a measure of estimation variability. The maximum SEM values are: less than 0.014 in Figure 4(a), less than 0.018 in Figure 4(b) and less than 0.015 in Figure 4(c). In the legend, “adapt” refers to the adapted COT value estimator, “ridge” refers to the ridge regression–based DualBounds method, and “knn” refers to the KNN–based DualBounds method.
In this section, we compare the estimation accuracy of the adapted COT value estimator proposed in Section 4.2 and the DualBounds method under the same three settings as described in Section 4.2, except that the propensity score is non-constant, defined as
$\mathbb{P}(W=1|Z=z)=e(z)=\frac{1}{1+e^{-\frac{3z}{2}}},\quad z\in\mathbb{R}.$
To investigate the effect of reweighting on COT value estimation (without being affected by errors in the propensity score estimator), we experiment with the scenario where the true function $e(\cdot)$ is fully known to all methods under evaluation. The results are presented in Figure 4.
From the empirical results, we observe the following: (a) In the linear location model (Figure 4(a)), the ridge-based DualBounds method outperforms the other approaches, owing to its use of a linear nuisance estimator that aligns well with the true model structure. (b) In both the quadratic location model (Figure 4(b)) and the scale model (Figure 4(c)), the adapted COT value estimator significantly outperforms the other two methods. (c) Compared to the results in Figure 2 (Section 4.2), the accuracy of estimation decreases in both the linear and quadratic location models, whereas the performance in the scale model remains relatively unchanged. One possible explanation is, in the scale model, the sign of $f_{k}(Z)~{}k=0,1$ is independent of the outcome. As a result, the correlation between $Y$ and $Z$ is weaker than in the other two models, reducing the influence of the reweighting induced by the propensity score.
5.2 Real Data

### v2

@@ 4.2 Covariate-Dependent Treatment
Unlike Bernoulli design where treatment assignment is independent of covariates by design, observational studies typically exhibit covariate-dependent treatment assignment. This results in non-constant propensity scores $e(z)=\mathbb{P}(W_{i}=1\mid Z_{i}=z),$ reflecting heterogeneity in treatment likelihood across covariate profiles. In this section, we discuss how the adapted COT value estimator can be tailored to this setting.
@@
Under the unconfoundedness assumption (Assp. 2), the conditional distributions of potential outcomes given covariates remain identifiable, i.e., consistent with $\mu_{Y(w)}^{z}$ for $w=0,1$. However, since the treatment assignment probability depends on covariates, the marginal distribution of covariates may differ between the treatment group and the control group, inducing a covariate shift. To formalize this setting with covariate shift, we consider the following super-population model of $Y(0),Y(1),Z,W$ following the joint distribution $\mu=\mu_{Y(0),Y(1),Z,W}$.
Assumption 10 (Covariate shift).
Assume that $((Z_{i}(w),Y_{i}(w)),i\in[n_{w}])$ are drawn i.i.d. from $\mu_{Z|W=w}\otimes\mu_{Y(w)}^{Z}$ for $w=0,1$.
@@
Since our estimand of interest $V^{*}$ relies on the marginal covariate distribution, irrespective of treatment assignment, it is essential to correct for the distributional shift induced by non-constant propensity scores. Specifically, we reweight the samples to align with the marginal covariate distribution. By Proposition 1, when the propensity score $e(z)$ is known, the appropriate population-level weighting is given by
$\xi_{w,i}=\frac{\mu_{Z}(Z_{i}(w))}{\mu_{Z\mid W=w}(Z_{i}(w))}\propto\frac{1}{1-w-e(Z_{i}(w))},~W_{i}=w.$
@@
On finite samples, when the propensity score $e(z)$ is unknown but estimated by $\hat{e}(z)$, we adopt a group-wise self-normalized weight: for $w=0,1$, $\hat{\xi}_{w,G}$ is defined as
$\frac{(1-w-\hat{e}(\varphi_{r}^{n_{w}}(G)))^{-1}}{\sum_{G}(1-w-\hat{e}(\varphi_{r}^{n_{w}}(G)))^{-1}|\{i:Z_{i}(w)\in G\}|}$ (7)
@@
to ensure the weights sum to one. Here, $G\in\Phi_{r}^{n}$ is a cell corresponding to projection map $\varphi_{r}^{n}$, where $\Phi_{r}^{n}\triangleq\left\{(\varphi_{r}^{n})^{-1}(\{x\}):x\in\varphi_{r}^{n}([0,1]^{d_{Z}})\right\}$, and $\varphi_{r}^{n}(G)$ is the cell center of $G\in\Phi_{r}^{n}$. Note that the weight (7) is different from the common self-normalized weights that are based on the sample value of $Z$. This is used to adapt the construction of the adapted COT value estimator with discretization to the setting in Assp. 10. Finally, our reweighted adapted empirical distributions are defined by: for $w=0,1$,
$\hat{\bm{\mu}}_{Y(w),Z(w),n_{w}}$ $=\sum_{G\in\Phi_{r}^{n_{w}}}\sum_{i:Z_{i}(w)\in G}\hat{\xi}_{w,G}\delta_{Y_{i}(w),\varphi_{r}^{n_{w}}(Z_{i}(w))}.$
@@
We then still employ $\hat{\bm{\mu}}_{Y(w),Z(w),n_{w}},w=0,1$ along with $\hat{\pi}_{n_{0},n_{1}}$ defined in Eq. (6), to construct the adapted COT value estimator as formulated in Definition 4.
@@
Specifically, we implement a two-fold cross-fitting procedure to ensure the independence between the estimated propensity score $\hat{e}$ and the adapted empirical distributions $\hat{\bm{\mu}}$. We randomly split the data into two folds of approximately equal size. For each fold, we construct $\hat{e}$ using data from the other fold, and then construct $\hat{\bm{\mu}}_{Y(w),Z(w),n_{w}}$ on the current fold with the estimated propensity score plugged in. This yields two estimates of the lower bound, and we further average them to form our final estimator $\hat{V}_{n_{0},n_{1}}$.
[…]
Assumption 11 (Lipschitz $\hat{e}$).
@@
Assumption 13 (Average error of propensity score).
Assume that for $n_{0},n_{1}\geq 1$,
$\mathbb{E}\left[\frac{1}{n_{w}}\sum_{i=1}^{n_{w}}\left|n_{w}\hat{\xi}_{w,i}-\frac{{\rm d}\mu}{{\rm d}\mu_{|W=w}}(Z_{i}(w))\right|\right]$ $\leq C_{\textup{w}}n_{w}^{-r},$
where $C_{\textup{w}}$ is a constant independent of $n_{0},n_{1}$ and
$\hat{\xi}_{w,i}\triangleq\frac{(1-w-\hat{e}(Z_{i}(w)))^{-1}}{\sum_{l=1}^{n}(1-w-\hat{e}(Z_{l}(w)))^{-1}},w=0,1.$
[…]
Theorem 6 (Finite-sample complexity with covariate shift).
Under Assp. 3-5, 7-13 with $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, then
$\mathbb{E}\left[|\hat{V}_{n_{0},n_{1}}-V_{\textup{c}}|\right]\leq C^{\dagger}\gamma_{d_{Y},d_{Z}}(n_{0}\wedge n_{1}),$
where $C^{\dagger}$ is a constant that depends on $d_{Z},d_{Y}$, $\delta$ (Assp. 3), $L_{h}$ (Assp. 7), $L_{Z}$ (Assp. 8), $L_{e}$ (Assp. 11), $\eta$ (Assp. 12) and $C_{\textup{w}}$ (Assp. 13).
@@ 5 Simulation
5.1 Selection of Cell Size
In Definition 2, we design the discretization cells with edge length growing in the order of $n^{-r}$, resulting in $O(n^{rd_{Z}})$ small cells, in order to balance the bias-variance tradeoff with respect to the sample size $n$. In practice, the constant in $O(n^{rd_{Z}})$ may also affect the estimation accuracy. Specifically, in this section we set the number of cells to the nearest integer of $cn^{rd_{Z}}$.
We conduct a sensitivity analysis of our method with respect to $c$ under the quadratic location model (see Section 5.2 (b)) with $n_{0}=n_{1}=300$. As shown in Table 1, we compute the average relative estimation error (and their standard error), i.e., $\mathbb{E}[|\hat{V}_{n_{0},n_{1}}-V_{\textup{c}}|]/V_{\textup{c}}$ for $5$ choices of $c$, where $\mathbb{E}$ is approximated through $500$ Monte Carlo simulation runs. Compared to the baseline $c=1$, we can see that when $c$ is relatively small, increasing $c$ improves the estimation accuracy while its performance tends to stabilize as $c$ grows larger.
Table 1: Sensitivity to $c$ (baseline $c=1$).
$c$ Relative error $\Delta$ vs. baseline
0.6 0.21 (5.4e-3) $38.1\%$
0.8 0.17 (5.2e-3) $13.4\%$
1.0 0.15 (4.9e-3) 0.0
1.2 0.14 (4.8e-3) $6.4\%$
1.4 0.14 (4.9e-3) $7.8\%$
In practice, since the true value of $V_{\textup{c}}$ is unknown, we suggest the following approach to select a $c$ that achieves a relatively higher estimation accuracy: (i) given a dataset with size $N$, generate $B=50$ bootstrap samples, each with size $N$, (ii) compute the average $\hat{V}_{n_{0},n_{1}}$ value of the bootstrap samples for each candidate of $c$, and plot a curve on the average values, (iii) pick the “elbow” point of the curve (the point with largest distance to the line connecting the start and end point of the curve), and choose the corresponding $c$ value. An illustration is shown in Figure 1.
Figure 1: Selection of the elbow point.
(a) Linear location model
(Bernoulli design)
(b) Quadratic location model
(Bernoulli design)
(c) Scale model
(Bernoulli design)
(d) Linear location model
(Covariate-dependent)
(e) Quadratic location model
(Covariate-dependent)
(f) Scale model
(Covariate-dependent)
Figure 2: Estimation accuracy comparison between the adapted COT estimator and DualBounds. In the legend, “adapt” refers to the adapted COT estimator, “ridge” refers to the ridge regression–based DualBounds, and “knn” refers to the KNN–based DualBounds. The average relative error is computed over 500 Monte Carlo repetitions. For the uncertainty, we compute the standard error of the mean (SEM), i.e. $\textup{SD}/\sqrt{500}$, where SD is the standard deviation. The numerical values of SEM satisfy: (a): <0.005<0.005; (b)(c): <0.02<0.02; (d)(e)(f): <0.018<0.018.
Table 2: Average relative error (SEM) over 500 Monte Carlo repetitions on the STAR-based semi-synthetic data.
Method ($N=n_{0}+n_{1}=2n_{0}$) $N=400$ $N=800$ $N=1200$ $N=1600$
ridge regression–based DualBounds $4.0\text{e}{-2}$ (1.3e-3) $3.2\text{e}{-2}$ (1.0e-3) $2.9\text{e}{-2}$ (0.9e-3) $2.7\text{e}{-2}$ (0.8e-3)
KNN–based DualBounds $5.1\text{e}{-2}$ (1.6e-3) $4.5\text{e}{-2}$ (1.3e-3) $4.5\text{e}{-2}$ (1.1e-3) $4.4\text{e}{-2}$ (1.0e-3)
Our method $\mathbf{3.6e\text{-}2~(1.2e\text{-}3)}$ $\mathbf{2.9e\text{-}2~(1.0e\text{-}3)}$ $\mathbf{2.6e\text{-}2~(0.8e\text{-}3)}$ $\mathbf{2.5e\text{-}2~(0.8e\text{-}3)}$
5.2 Simulation Results
Synthetic data. To evaluate the effectiveness of the adapted COT value estimator in simulation [^3] The code and additional experimental results are available at github.com/siruilin1998/causalPIviaCOT., we introduce the following synthetic data-generating mechanism. For $w=0,1$, $Y(w)=F_{m}(f_{w}(Z),\varepsilon_{w})$, where $Z,\varepsilon_{w},Y(w)$ are independent with each other. The model $F_{m}$ includes: (i) location model $F_{1}(u,v)=u+v$, and (ii) scale model $F_{2}(u,v)=(u_{j}v_{j},j\in[d_{Y}])$. More concretely, we consider $d_{Y}=d_{Z}=1$, $Z,\varepsilon_{w}\sim N(0,1)$, and three models:
@@ 5.2 Simulation Results
For these Gaussian models, the $V_{\textup{c}}$’s population value has a closed form without referring to an optimization problem (see appendix for a proof), and is computable through a simple Monte Carlo algorithm with parametric convergence rate. Therefore, these models enable the evaluation of the estimation accuracy.
@@
We compare the estimation accuracy of our adapted COT value estimator with the method in [28], implemented in their Python package DualBounds [^4] https://dualbounds.readthedocs.io. Their approach relies on estimating nuisance functions associated with the covariate–outcome model. In our comparison, we evaluate both the ridge regression-based and $k$-nearest neighbor (KNN)-based DualBounds methods. Our implementation is based on the Python Optimal Transport (POT [^5] https://pythonot.github.io/) library [22].
@@
In Figure 2, we present the comparison in the setting of: (i) Bernoulli design ((a)(b)(c)) and (ii) Covariate-dependent treatment ((d)(e)(f)) with $e(z)=(1+e^{-\frac{3z}{2}})^{-1}\ \forall z\in\mathbb{R}$, which we assume to be fully known for all methods under evaluation. The figure shows that: (i) For the linear location model, the ridge-based DualBounds method performs best due to its use of a linear nuisance estimator that aligns with the true model structure. Nonetheless, the adapted COT value estimator achieves comparable accuracy despite not relying on model-specific assumptions (see (a)). (ii) For the quadratic location model, the adapted COT value estimator outperforms both versions of DualBounds and the KNN-based outperforms the ridge-based, as the latter cannot capture the nonlinearity in the true model. (iii) For the scale model, the adapted COT estimator also has the best performance.
Real data. We also consider the real-world dataset from the Student Achievement and Retention (STAR) Demonstration Project [1], where the academic performance (outcome) is measured by the GPA recorded in the first academic year and the baseline GPA serves as the covariate. To make the COT value accessible at the population level, we evaluate the methods using hypothetical data generated from a model fitted to the real data. Table 2 presents the accuracy on estimating the PI bounds of the correlation between two potential outcomes, and our adapted COT value estimator consistently outperforms the others.
6 Conclusion and Discussion
Conclusion.
In this paper, we study the optimal covariate-assisted partial identification sets for causal estimands by solving COT problems. Our finite-sample COT estimator avoids nuisance function estimation, does not require well-specified models, and enjoys a provable statistical convergence rate. Furthermore, our adapted COT value estimator answers the question proposed in the discussion of [35] on how to combine the adapted Wasserstein distance into the COT value estimation. Specifically, our adapted estimator is their $V_{\text{causal}}(\eta)$ when $\eta$ approached infinity.
Triangular transport maps.
We next discuss the connection between our COT framework and triangular transport maps. When the transport cost is quadratic and the conditional outcome distributions admit densities, the optimal coupling can be represented by a triangular transport map [14, 13, 26]. Such maps arise from a sequential alignment of conditional distributions and coincide, in the one-dimensional outcome case, with the classical Knothe–Rosenblatt rearrangement [43, 33].
Triangular transport has been used in several causal modeling and inference frameworks [15, 19, 6], where it is often interpreted as a structural mechanism for generating counterfactual outcomes. Our perspective is complementary. Rather than modeling unit-level counterfactuals via an estimated transport map, we use the induced causal coupling structure to define and estimate a distributional functional. In particular, our estimator targets the COT value directly, without requiring explicit construction of the underlying triangular map.
Nevertheless, we conjecture that our discretization approach (Section 4.1.1) could be adapted to recover a triangular transport map in a manner analogous to plug-in constructions in the unconditional setting; see, for example, the one-nearest-neighbor estimator in [36, Section 4.2]. In that case, the estimation error of the induced triangular transport gap would be expected to match the order of the estimation error for the COT value itself.
Multiple treatment levels.
We discuss here the possibilities of an extension to settings with multiple treatment levels. Suppose the treatment variable takes more than two levels. A natural extension of our framework is to consider multi-marginal optimal transport, which constructs a joint coupling across all conditional outcome distributions given the covariates. This provides a geometric framework for jointly aligning multiple potential outcome distributions. The unconditional multi-treatment setting has been studied in [24].
At the same time, the use of transport maps in counterfactual modeling has important limitations. Recent work [17, 18] shows that optimal transport maps generally cannot serve as structural models for counterfactual assignments when more than two treatment levels are present. We emphasize that our methodology does not interpret OT maps as structural counterfactual mechanisms. Instead, OT and COT are employed as distributional comparison tools: they measure discrepancies between conditional outcome laws and construct PI intervals via geometric constraints on admissible couplings. Consequently, conditional OT remains meaningful in multi-treatment settings for constructing PI intervals and bounding functionals of potential outcomes.
Acknowledgements
The material in this paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Additional support is gratefully acknowledged from NSF 2118199, 2229012, 2312204, 2403007, and ONR 13983111.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Yes]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Yes]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Yes]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Yes]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Yes]
Supplementary Materials
Organization. The appendix is organized as follows. Section A presents the details of the real data experiment discussed in Section 5.2. Section B presents supplementary experiments to show the robustness and effectiveness of our method. Section C reviews technical background knowledge. Section D provides proofs for the results in Section 2. Section E provides proofs for the results in Section 3. Section F provides proofs for the results in Section 4. Section H collects proofs for supporting lemmas.
Appendix A More Details on the Real Data Experiment